# ShapeWords: Guiding Text-to-Image Synthesis with 3D Shape-Aware Prompts

### v1

@@ 5 Conclusion
Our supplement shows failure cases for our method. In summary, it struggles with capturing challenging fine-grained geometry, especially in areas of thin parts. Generalization to much more complex prompts would also requiring training on much larger scale datasets. Finally, our method exhibits certain appearance biases (e.g. oversaturation of some images, or preference of certain colors), which could be avoided with better score distillation variants. Combining our shape tokens with viewpoint and other 3D attribute tokens would also be an interesting future direction [7, 9].
@@ 6.1 Data generation details
For depth images, we used the inverted ShapeNet data provided by the ULIP authors [47]. As stated in the main paper, for each depth image, we applied a randomly selected prompt from a set of $13,716$ prompts for ControlNet conditioning. The ControlNet-based generation was done for $50$ steps with control strength of $2$. To promote adherence to shape geometry and reduce appearance biases during training, we additionally used the Stable Diffusion 2.1 inpainting model [38] for 50 steps to modify the backgrounds while preserving the foreground objects. The inpainting strength was set to $0.5$.
@@ 6.2 SDS weighting function
For the SDS optimization loss of Eq. 6, we use the weighting function $W(t)$, proposed by DreamTime [18], that enhances training stability:
@@ 7 Running times
We note that ShapeWords and ControlNet-based baselines rely on the same Stable Diffusion model (Stable Diffusion 2.1 base) and have similar computation requirements at test time: given a text prompt or/and depth image, it takes a few seconds to generate an image with $100$ diffusion steps on a single GPU: $6.79$s for ControlNet; and $5.00$s for Stable Diffusion 2.1 with ShapeWords. The forward pass of the Shape2CLIP module takes $0.003$s with pre-computed PointBERT embeddings. All measurements were done on single A40 GPU with batch size $1$ and based on an average across 20 runs. For details on PointBERT computational costs, we refer to [49].

### v2

@@ 5 Conclusion
Our supplement shows failure cases for our method. In summary, it struggles with capturing challenging fine-grained geometry, especially in areas of thin parts. Generalization to much more complex prompts would also requiring training on much larger scale datasets. Finally, our method exhibits certain appearance biases (e.g. oversaturation of some images, or preference of certain colors), which could be avoided with better score distillation variants. Combining our shape tokens with viewpoint and other 3D attribute tokens would also be an interesting future direction [viewneti, cheng2024learning].
@@ 6.1 Data generation details
For depth images, we used the inverted ShapeNet data provided by the ULIP authors [xue2023ulip]. As stated in the main paper, for each depth image, we applied a randomly selected prompt from a set of $13,716$ prompts for ControlNet conditioning. The ControlNet-based generation was done for $50$ steps with control strength of $2$. To promote adherence to shape geometry and reduce appearance biases during training, we additionally used the Stable Diffusion 2.1 inpainting model [rombach2022stable] for 50 steps to modify the backgrounds while preserving the foreground objects. The inpainting strength was set to $0.5$.
@@ 6.2 SDS weighting function
For the SDS optimization loss of Eq. 6, we use the weighting function $W(t)$, proposed by DreamTime [huang2024dreamtime], that enhances training stability:
@@ 7 Running times
We note that ShapeWords and ControlNet-based baselines rely on the same Stable Diffusion model (Stable Diffusion 2.1 base) and have similar computation requirements at test time: given a text prompt or/and depth image, it takes a few seconds to generate an image with $100$ diffusion steps on a single GPU: $6.79$s for ControlNet; and $5.00$s for Stable Diffusion 2.1 with ShapeWords. The forward pass of the Shape2CLIP module takes $0.003$s with pre-computed PointBERT embeddings. All measurements were done on single A40 GPU with batch size $1$ and based on an average across 20 runs. For details on PointBERT computational costs, we refer to [yu2022pointbert].